Evaluate the integral of (x³- 2x+5)/(x-3) dx


1
Expert's answer
2021-03-31T08:47:23-0400

Convert the expression "\frac{x^3-2x+5}{x-3}" so that the numerator contains a polynomial of smaller degree than the denominator

"\frac{x^3-2x+5}{x-3}=\frac{x^3-3x^2+3x^2-9x+7x-21+26}{x-3}="


"=\frac{x^2(x-3)+3x(x-3)+7(x-3)+26}{x-3}=x^2+3x+7+\frac{26}{x-3}"

Therefore, we can start evaluating the integral,


"\intop\frac{x^3-2x+5}{x-3}dx=\intop(x^2+3x+7+\frac{26}{x-3})dx="

"=\frac{x^3}{3}+\frac{3x^2}{2}+7x+26 ln|x-3|+C"